# Subsystem many-hypercube codes: High-rate concatenated codes with low-weight syndrome measurements

### v1

@@ Abstract
Quantum error-correcting codes (QECCs) require high encoding rate in addition to high threshold unless a sufficiently large number of physical qubits are available. The many-hypercube (MHC) codes defined as the concatenation of the [[6,4,2]][\![6,4,2]\!] quantum error-detecting code have been proposed as high-performance and high-encoding-rate QECCs. However, the concatenated codes have a disadvantage that the syndrome weight grows exponentially with respect to the concatenation level. To address this issue, here we propose subsystem quantum codes based on the MHC codes. In particular, we study the smallest subsystem MHC codes, namely, subsystem codes derived from the concatenated [[4,2,2]][\![4,2,2]\!] error-detecting codes. The resulting codes have a constant syndrome-measurement weight of 4, while keeping high encoding rates. We develop the block-MAP and neural-network decoders and show that they demonstrate superior performance to the bounded-distance decoder.
@@ I Introduction
In this work, we propose subsystem codes made from the MHC codes [8], which are referred to as the subsystem MHC codes. As a simple example, we study the [[4,2,2]][\![4,2,2]\!] subsystem MHC codes, which are the subsystem codes of the concatenated [[4,2,2]][\![4,2,2]\!] error-detecting codes. The [[4,2,2]][\![4,2,2]\!] code is the smallest stabilizer code that can detect a single-qubit error [27, 28]. The resulting code at concatenation level $r$ is a $4^{r},2^{r},g,2^{r}]\!]$ code, where $g=4^{r}+2^{r}-2\cdot 3^{r}$ is the number of gauge qubits. The syndrome measurements are performed on four qubits at an arbitrary level, while the original level-$r$ [[4,2,2]][\![4,2,2]\!] MHC code has check operators of weight as high as $4\cdot 2^{r-1}$. However, the decoding becomes problematic due to a large number of gauge qubits. We overcome this problem by developing two kinds of decoders for the new codes: a block MAP (maximum-a-posteriori) decoder [29] up to level 3 and a neural-network (NN) decoder [30, 31] up to level 4. We evaluate the performance of the [[4,2,2]][\![4,2,2]\!] subsystem MHC codes and show that they outperform the theoretical performance of the bounded-distance (BD) decoder and also the performance evaluated with the belief propagation with ordered statistics decoding (BP-OSD) [32].
@@ II.2 $D_{4,4}$ code
Figure 2: The operators of (a) the $D_{4,4}$ code and (b) the subsystem $D_{4,4}$ code. The logical-qubit operators are the same for the two codes. (c) Operators of the subsystem $D_{4,4,4}$ code. Note that only some of the stabilizer generators are depicted.
@@ II.4 Subsystem $D_{4^{r}}$ code
II.5 Encoding and syndrome measurement
@@ III Decoders
In the following, we evaluate the performance of the subsystem $D_{4^{r}}$ codes up to $r=4$ assuming bitflip errors. The other errors such as measurement errors or circuit-level noise are not considered and left for future work. We used Stim [35] to simulate the syndrome outcome.
Our codes have a large number of gauge qubits at higher levels, that is, less information for decoding. This makes some elaborate decoders infeasible at least within reasonable calculation time. For this reason, we develop a NN decoder to empirically learn the decoding up to level 4. In addition, we use the block MAP decoder, which is optimal, up to level 3.
@@ IV Conclusions
Due to the presence of exponentially growing number of gauge qubits, the numerical cost of exhaustive searches, such as the brute-force or block MAP decoding, grows far faster than the original [[4,2,2]][\![4,2,2]\!] MHC codes as the concatenation level increases. Thus, we have developed a fully-connected deep NN decoder to reduce the numerical cost in decoding. The performance of the NN decoder exceeds that of the BD decoder and the BP-OSD. Compared with the original [[4,2,2]][\![4,2,2]\!] MHC codes, the estimated threshold for the bitflip error is reduced from 7% to 2% (see Appendix A). This reduction could be due to reduced information from the syndrome measurements, and also due to the decoder that is not sufficiently optimized. Indeed, the performance of the NN decoder has not reached that of the block MAP (optimal) decoder. Closing this gap would be a potential future work.

### v2

@@ Abstract
Quantum error-correcting codes (QECCs) require high encoding rate in addition to high threshold unless a sufficiently large number of physical qubits are available. The many-hypercube (MHC) codes defined as the concatenation of the [[6,4,2]][\![6,4,2]\!] quantum error-detecting code have been proposed as high-performance and high-encoding-rate QECCs. However, the concatenated codes have a disadvantage that the syndrome weight grows exponentially with respect to the concatenation level. To address this issue, here we propose subsystem quantum codes based on the MHC codes. In particular, we study the smallest subsystem MHC codes, namely, subsystem codes derived from the concatenated [[4,2,2]][\![4,2,2]\!] error-detecting codes. The resulting codes have a constant syndrome-measurement weight of 4, while keeping high encoding rates. We build the block-MAP and neural-network decoders and show that they demonstrate superior performance to the bounded-distance decoder.
@@ I Introduction
In this work, we propose subsystem codes made from the MHC codes [8], which are referred to as the subsystem MHC codes. As a simple example, we study the [[4,2,2]][\![4,2,2]\!] subsystem MHC codes, which are the subsystem codes of the concatenated [[4,2,2]][\![4,2,2]\!] error-detecting codes. The [[4,2,2]][\![4,2,2]\!] code is the smallest stabilizer code that can detect a single-qubit error [27, 28]. The resulting code at concatenation level $r$ is a $4^{r},2^{r},g,2^{r}]\!]$ code, where $g=4^{r}+2^{r}-2\cdot 3^{r}$ is the number of gauge qubits. The syndrome measurements are performed on four qubits at an arbitrary level, while the original level-$r$ [[4,2,2]][\![4,2,2]\!] MHC code has check operators of weight as high as $4\cdot 2^{r-1}$. However, the decoding becomes problematic due to a large number of gauge qubits. We overcome this problem by building two kinds of decoders for the new codes: a block MAP (maximum-a-posteriori) decoder [29] up to level 3 and a neural-network (NN) decoder [30, 31] up to level 4. We evaluate the performance of the [[4,2,2]][\![4,2,2]\!] subsystem MHC codes and show that they outperform the theoretical performance of the bounded-distance (BD) decoder and also the performance evaluated with the belief propagation with ordered statistics decoding (BP-OSD) [32].
@@ II.1 $D_{4}$ code
This syndrome extraction method can detect harmful errors that spread into data qubits [34].
@@ II.2 $D_{4,4}$ code
Figure 2: The logical-qubit operators, gauge-qubit operators, stabilizer generators and check operators of (a) the $D_{4,4}$, (b) subsystem $D_{4,4}$, (c) subsystem $D_{4,4,4}$, (d) $D_{6,6}$, (e) subsystem $D_{6,6}$, and (f) subsystem $D_{6,6,6}$ codes are shown. In (c) and (f), the stabilizer generators that extend along the $x$, $y$, and $z$ directions are shown separately. The logical-qubit (gauge-qubit) operators are defined on qubits shown by red squares or cubes (yellow lines). Notice that the logical $X$ and $Z$ operators of the same logical or gauge qubits are not defined on the same sets of qubits; they share an odd number of qubits so that they anticommute with each other. Two types of stabilizer generators $SX$ and $SZ$ are defined on the same sets of qubits shown by blue lines, rectangles or cuboids. This is also the case for the $X$ and $Z$ check operators (black lines).
@@ II.4 Subsystem $D_{4^{r}}$ codes
II.5 Encoding and syndrome measurements of $D_{4^{r}}$
@@
II.6 Subsystem $D_{6^{r}}$ codes
The construction of the subsystem $D_{6^{r}}$ codes from the original $D_{6^{r}}$ codes is done in a similar way to that of the subsystem $D_{4^{r}}$ codes. Compared with the original $D_{6^{r}}$ codes, the syndrome-measurement weight is reduced from $6\cdot 2^{r-1}$ to 6.
The $D_{6^{r}}$ codes are defined on qubits placed on an $r$-dimensional hyperlattice of linear dimension 5. There are $r\cdot 6^{r-1}$ lines of length 5. We define weight-6 check operators as the product of $X$ and $Z$ on these lines, which form the gauge group. The procedure of constructing the subsystem $D_{6^{r}}$ codes from this point onward is similar to that of the subsystem $D_{4^{r}}$ codes in Sec. II.4. The resulting code has $4^{r}$ logical qubits, $2(5^{r}-4^{r})$ stabilizer generators, and $6^{r}+4^{r}-2\cdot 5^{r}$ gauge qubits. Specifically, the operators of the $D_{6,6}$, subsystem $D_{6,6}$, and subsystem $D_{6,6,6}$ codes are shown schematically in Figs. 2 (d)–2 (f). We assign gauge-qubits operators within the hyperplanes at $x_{n}=4\,(n\in[1,r])$.
@@ III Decoders
In the following, we focus on the subsystem $D_{4^{r}}$ codes and evaluate their performance up to $r=4$ assuming bitflip errors. The other errors such as measurement errors or circuit-level noise are not considered and left for future work. We used Stim [35] to simulate the syndrome outcome.
Our codes have a large number of gauge qubits, the number of which grows exponentially with respect to the concatenation level. Compared with the original $D_{4^{r}}$ codes, the number of the stabilizer generators decreases by that of the gauge qubits, that is, less information are available for decoding. In addition, the number of error degeneracy is $2^{\text{\# of the gauge qubits}}$ as large as the $D_{4^{r}}$ codes since any two errors that are equivalent modulo gauge-qubit operators are regarded as the same error. This makes optimal or nearly optimal decoders, which enumerate the all error degeneracy, infeasible at least within reasonable calculation time. For this reason, we build a NN decoder to empirically learn the decoding up to level 4. In addition, we use the block MAP decoder, which is optimal, up to level 3.
@@ IV Conclusions
Due to the presence of exponentially growing number of gauge qubits, the numerical cost of exhaustive searches, such as the brute-force or block MAP decoding, grows far faster than the original [[4,2,2]][\![4,2,2]\!] MHC codes as the concatenation level increases. Thus, we have devd a fully-connected deep NN decoder to reduce the numerical cost in decoding. The performance of the NN decoder exceeds that of the BD decoder and the BP-OSD. Compared with the original [[4,2,2]][\![4,2,2]\!] MHC codes, the estimated threshold for the bitflip error is reduced from 7% to 2% (see Appendix A). This reduction could be due to reduced information from the syndrome measurements, and also due to the decoder that is not sufficiently optimized. Indeed, the performance of the NN decoder has not reached that of the block MAP (optimal) decoder. Closing this gap would be a potential future work.
Compared with the teleportation-based error correction proposed in [8], the advantages of the proposed syndrome extraction by ancillae measurements are that the number of ancillae is reduced by a factor of 1/4 (1/6) for the subsystem $D_{4^{r}}$ ($D_{6^{r}}$) codes, and that this method is free from preparing logical Bell states. However, the syndrome extraction with ancillae in general requires repeated measurements, which would make the time overhead larger than the teleportation-based one. In addition, entangling data qubits with ancillae could lead to error spreading, which is not considered in this work. We will discuss the circuit-level noise separately in the forthcoming paper.